A capacitor is completely charged with 600 nC by a voltage source that had 325 V.

What is its capacitance?

The initial air gap of the capacitor above was 5 mm. What is the stored energy if the air gap is now 13 mm?
Answer #1

capacitance C = charge/ potential = 600/325 = 1.846 nF

capacitance is inversely proportional to air gap.

so new capacitance = 1.846*5/13 =0.71 nF

hence new energy = q2 /2*C = 6002 /2*0.71 = 0.253 milli farad